# supplement

### v2

@@ 7.4 Demonstration of acceptable truncation error
In this section, we demonstrate the upper bounds for truncation error in (25) and Theorem 5.2, Theorem 5.5 required to ensure truncated subspace iteration converges. To this end, we obtain the eigenvectors and eigenvalues with standard methods and use them to compute the upper bounds for acceptable truncation error at each iteration. We considered the Laplacian (53) with $d=2$, $n=32$ and used a Chebyshev polynomial filter of degree $k=8$, resulting in an eigenvalue ratio $|p_{k}(\lambda_{2})/p_{k}(\lambda_{1})|\approx 0.6807$. At each iteration, we computed the upper bound in (25) for the acceptable truncation error and adaptively selected the truncation rank with the TT-SVD truncation algorithm so that the inequality is satisfied. Fig. 4(a) shows both the computed upper bound for the truncation error and the actual truncation error achieved using the rank selected by the TT-SVD truncation algorithm. Fig. 4(b) shows the rank selected by the TT-SVD truncation at each iteration. In early iterations, achieving the small truncation errors demands a larger rank. This is because applying the operator $p_{k}(A)$ to a random initial vector generates a vector that require higher ranks for an accurate representation. As the algorithm nears convergence, small truncation errors are again required; however, this can be achieved with a smaller rank since the dominant eigenvector $\psi_{1}$ has rank $1$. Finally, Fig. 4(c) shows the angle $\theta$ of the truncated power iteration vector with the dominant eigenvector $\psi_{1}$, demonstrating that the method is converging.
@@
Figure 5: Numerical demonstration of the truncation error upper bounds in (25) and Theorem 5.2 during the computation of the dominant eigenvector of the Laplacian (53) with dimension $d=2$ and $n=32$ points per dimension. A Chebyshev polynomial filter of degree $8$ was used, yielding an eigenvalue ratio $|p_{k}(\lambda_{2})/p_{k}(\lambda_{1})|\approx 0.6807$. The figure shows the projections of iterates onto $\mathbb{R}^{2}$, as defined in (23), at various iterations both before and after truncation.
@@ 7.5 Methods for improving efficiency
8 4 0.20s 3106 621s $<10^{-10}$
8 8 0.44s 681 300s $<10^{-10}$
To investigate the convergence of the subspace iteration, we considered a Heisenberg spin-1/2 chain of length $L=32$ and computed the first excited state using subspaces of dimensions $m=2,4,8$ and polynomial filters of degrees $k=2,4,8$. Increasing either the subspace dimension or the polynomial filter degree results in more costly iterations. However, similar to classical subspace iteration, increasing these parameters accelerates convergence, demonstrating the robustness of subspace iteration to rank truncations. For all experiments we truncate with a maximum rank of $2$ during each iteration. Table 1 presents the average CPU-time per iteration, the number of iterations, and the residual of the first excited state.
Finally, we demonstrate two strategies for improving the computational efficiency of rank-truncated subspace iteration. First, using randomized algorithms to compute low-rank approximations of linear combinations in the Rayleigh-Ritz step as described in Section 6.2. Second, approximating matrix-vector products and linear combinations in appropriate tangent spaces of low-rank TT manifolds as described in Section 2.1.4.
Figure 6: Comparison between TT-SVD and randomized truncation methods [4] for approximating linear combinations in the Rayleigh-Ritz step of subspace iteration. We considered the spin-1/2 Heisenberg chain of length $L=24$, iterating at fixed TT-rank $r=16$. For both methods, TT-SVD is used for truncations during polynomial filtering. (a) Comparison of cpu time. Using of the randomized truncation methods consistently reduces the per-iteration run time of the Rayleigh-Ritz step by more than 2.5x. This speed up is particular important for larger subspaces, for which the Rayleigh-Ritz step requires increasingly more time than polynomial filtering. (b) Comparison of convergence. number of iterations required for each eigenvector estimate to converge in residual $\left\lVert Av_{j}/\langle v_{j},Av_{j}\rangle-v_{j}\right\rVert_{F}<10^{-12}$ for a subspace of size $m=96$. We plot the min, mean, and max over 6 different initial subspaces. Despite the potential for additional error due to randomization, using randomized methods for truncation during the Rayleigh Ritz step prolongs convergence by no more than 3 iterations on average.
@@
To demonstrate the computational advantage of approximating matrix-vector products and linear combinations using orthogonal projection onto TT tangent spaces as described in Section 2.1.4 we considered the Hamiltonian (54) as a TT matrix (4) with accuracy $10^{-10}$ yielding TT matrix rank $\bm{r}_{A}=\begin{bmatrix}1&4&4&4&4&1\end{bmatrix}$. The subspace dimension was set to $m=15$, and a Chebyshev polynomial filter of $k=6$ was used. The method was run with maximum truncation rank $r_{\textrm{max}}=32$ and different rank truncation strategies. First we used TT-SVD truncation to approximate matrix-vector products and linear combinations of TTs. Such approach is computationally expensive since, e.g., matrix-vector products between $H$ and $v$ result in TTs with ranks equal to the product of the ranks of $H$ and $v$ which are then truncated costing $\mathcal{O}(dnr_{v}^{3}r_{A}^{3})$, where $r_{v}$ and $r_{A}$ are the entries of the corresponding TT-rank vectors (here assumed to be constant). Then we ran the subspace iteration and approximated matrix-vector products and linear combinations in appropriate tangent spaces (TS) of low-rank TT manifolds. This approach avoids rank increase and is therefore more efficient, e.g., with approximate matrix-vector products between $A$ and $v$ costing $\mathcal{O}(dnr_{v}^{3})$ where $Av$ is approximated in the tangent space of the low-rank manifold at $v$. Fig. 7 shows the norm of the residual for each Ritz vector at each iteration. As the maximum truncation rank increases, the final accuracy of the Ritz vectors improves. Notably, different truncation ranks and truncation strategies (SVD or TS) do not significantly affect the convergence rate of the method, even when they influence the final accuracy. These observations highlight the robustness of the subspace iteration convergence rate to rank truncation errors. The $100$ iterations using TT-SVD truncation took 1896 seconds while $100$ iterations with tangent space projections took 656 seconds, more than 2.5x speed-up.
@@ 8 Conclusions
We examined low-rank inexact variants of the Lanczos method and a polynomial filtered subspace iteration for tensor eigenvalue problems. These methods typically avoid the issue of getting stuck at local minimizers, which can occasionally occur in DMRG-based algorithms. Furthermore, these methods can be easily used to compute more than one eigenpair, whereas the original DMRG algorithm is designed to compute the lowest eigenvalue (i.e., the ground state). Our results demonstrated that orthogonal Krylov basis vectors often cannot be accurately represented by low-rank tensor trains, even when the eigenvectors of the associated matrix exhibit low-rank structure. We further showed that the low-rank approximation errors from representing orthogonal Krylov basis vectors resulted in stagnation of convergence in the Lanczos method. In contrast, the subspace iteration method employs Ritz vectors as the subspace basis, which directly approximate the low-rank eigenvectors of interest. We provided a convergence analysis quantifying the acceptable truncation error required to ensure progress toward convergence at each iteration. Numerical experiments further confirmed that the subspace iteration method exhibits significantly greater robustness to truncation errors compared to the Lanczos method, indicating it is a more reliable low-rank approach for computing several extreme eigenpairs of rank-structured matrices. We provided numerical demonstrations of the convergence analysis and showed that polynomial filtered subspace iteration can be used to compute several dominant eigenpairs with machine precision for problems with up to $100$ dimensions, provided the eigenvectors admit low-rank representations. We further demonstrated that rank-truncated subspace iteration can converge for problems where the density matrix renormalization group method (DMRG). Although the analysis presented here focuses on polynomial filtered subspace iteration, it can be easily extended to other type of subspace iteration such as the one based on the FEAST algorithm [2].
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program under U.S. Department of Energy Contract No. DE-AC02-05CH11231.
Appendix A Proof of Theorem 5.1
Proof A.1.
Tangent of the angles $\measuredangle(v,\psi_{1})$ and $\measuredangle(Av,\psi_{1})$ can be expressed as
(55) $\tan\measuredangle(v,\psi_{1})=\frac{\left\|\mathcal{P}_{\perp}v\right\|}{\left\|\mathcal{P}v\right\|}\quad\text{and}\quad\tan\measuredangle(Av,\psi_{1})=\frac{\left\|\mathcal{P}_{\perp}Av\right\|}{\left\|\mathcal{P}Av\right\|}.$
@@
(56) $\left\|\mathcal{P}_{\perp}Av\right\|\leq|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|\quad\text{and}\quad\left\|\mathcal{P}Av\right\|=|\lambda_{1}|\left\|\mathcal{P}v\right\|.$
@@
(57) $\left\|\mathcal{P}_{\perp}Av\right\|==\left\|\sum_{j=2}^{n_{1}\cdots n_{d}}\alpha_{j}\lambda_{j}\psi_{j}\right\|\\ \leq|\lambda_{2}|\left\|\sum_{j=2}^{n_{1}\cdots n_{d}}\alpha_{j}\psi_{j}\right\|\\ =|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|,$
@@
(58) $\tan\measuredangle(Av,\psi_{1})=\frac{\left\|\mathcal{P}_{\perp}Av\right\|}{\left\|\mathcal{P}Av\right\|}\leq\frac{|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|}{|\lambda_{1}|\left\|\mathcal{P}v\right\|}\\ =\left|\frac{\lambda_{2}}{\lambda_{1}}\right|\tan\measuredangle(v,\psi_{1}),$
completing the proof.
Appendix B Proof of Theorem 5.5
Proof B.1.
Using the eigendecomposition $A=\Psi\Lambda\Psi^{\text{T}}$, we write $T_{AV}$ as
(59) $T_{AV}$ $=\Psi_{>m}^{\text{T}}AV\left(\Psi_{\leq m}^{\text{T}}AV\right)^{-1}$
@@
(60) $\tan\Theta(\mathcal{R}(AV),\mathcal{R}(\Psi_{\leq m}))$ $=\sigma_{1}(T_{AV})$

### v3

@@ 6.4 Demonstration of acceptable truncation error
To demonstrate the upper bounds for truncation error in (18) and theorem 4.2, theorem 4.3 which ensure truncated subspace iteration converges, we considered the Laplacian (47) with $d=2$ and $n=32$. We used a Chebyshev polynomial filter of degree $k=8$, resulting in an eigenvalue ratio $|p_{k}(\lambda_{2})/p_{k}(\lambda_{1})|\approx 0.6807$. At each iteration we computed the upper bound in (18) for the acceptable truncation error and adaptively selected the rank using TT-SVD so that (18) is satisfied. fig. 4(a) shows the computed upper bound for the truncation error and the actual truncation error achieved using the rank selected by the TT-SVD truncation algorithm. fig. 4(b) shows the rank selected by the TT-SVD truncation at each iteration. In early iterations, achieving the small truncation errors demands a larger rank. This is because applying the operator $p_{k}(A)$ to a random initial vector generates a vector that require higher ranks for an accurate representation. As the algorithm converges, small truncation errors are again required; however, this can be achieved with a smaller rank since the dominant eigenvector $\psi_{1}$ has rank $1$. Finally, fig. 4(c) shows the angle $\theta$ of the truncated power iteration vector with the dominant eigenvector $\psi_{1}$, demonstrating that the method is converging.
@@
Figure 5: Numerical demonstration of the truncation error upper bounds in (18) and theorem 4.2 during the computation of the dominant eigenvector of the Laplacian (47) with dimension $d=2$ and $n=32$ points per dimension. A Chebyshev polynomial filter of degree 88 was used, yielding an eigenvalue ratio $|p_{k}(\lambda_{2})/p_{k}(\lambda_{1})|\approx 0.6807$. The figure shows the projections of iterates onto $\mathbb{R}^{2}$, as defined in (16), at various iterations both before and after truncation.
@@ 6.5 Methods for improving efficiency
8 4 0.20s 3106 621s $10^{-10}$
8 8 0.44s 681 300s $10^{-10}$
To compare the choice of subspace size and polynomial filter degree in rank-truncated subspace iteration, we considered a Heisenberg spin-1/2 chain of length $L=32$ and computed the first excited state using subspaces of dimensions $m=2,4,8$ and polynomial filters of degrees $k=2,4,8$. Increasing either the subspace dimension or the polynomial filter degree results in more costly iterations. However, similar to classical subspace iteration, increasing these parameters accelerates convergence. For all experiments we truncate with a maximum rank of $2$ during each iteration. table 1 presents the average CPU-time per iteration, the number of iterations, and the residual of the first excited state.
Finally, we compare implementations of the rank-truncation operator $\mathfrak{T}_{\bm{r}}$ for improving the computational efficiency of rank-truncated subspace iteration. First, we compare randomized algorithms with TT-SVD truncation to compute low-rank approximations of linear combinations in the Rayleigh-Ritz step as described in section 5.2. Second, we compare tangent space projections with TT-SVD for approximating matrix-vector products and linear combinations.
Figure 6: Comparison between $T$-SVD and randomized truncation methods [4] for approximating linear combinations in the Rayleigh-Ritz step of subspace iteration. We considered the spin-1/2 Heisenberg chain of length $L=24$, iterating at fixed $T$-rank $r=16$. For both methods, $T$-SVD is used for truncations during polynomial filtering. (a) Comparison of cpu time. Using of the randomized truncation methods consistently reduces the per-iteration run time of the Rayleigh-Ritz step by more than 2.5x. This speed up is particular important for larger subspaces, for which the Rayleigh-Ritz step requires increasingly more time than polynomial filtering. (b) Comparison of convergence. number of iterations required for each eigenvector estimate to converge in residual $\left\lVert Av_{j}/\langle v_{j},Av_{j}\rangle-v_{j}\right\rVert_{F}<10^{-12}$ for a subspace of size $m=96$. We plot the min, mean, and max over 6 different initial subspaces. Despite the potential for additional error due to randomization, using randomized methods for truncation during the Rayleigh Ritz step prolongs convergence by no more than 3 iterations on average.
@@
To demonstrate the computational advantage of approximating matrix-vector products and linear combinations using orthogonal projection onto TT tangent spaces as described in appendix B we considered the Hamiltonian (48) as a TT matrix (4) with accuracy $10^{-10}$ yielding TT matrix rank $\bm{r}_{A}=\begin{bmatrix}1&4&4&4&4&1\end{bmatrix}$. The subspace dimension was set to $m=15$, and a Chebyshev polynomial filter of $k=6$ was used. The method was run with maximum truncation rank $r_{\textrm{max}}=32$ and different rank truncation strategies. First we used TT-SVD truncation to approximate matrix-vector products and linear combinations of TTs. Such approach is computationally expensive since, e.g., matrix-vector products between $H$ and $v$ result in TTs with ranks equal to the product of the ranks of $H$ and $v$ which are then truncated costing $\mathcal{O}(dnr_{v}^{3}r_{A}^{3})$, where $r_{v}$ and $r_{A}$ are the entries of the corresponding TT-rank vectors (here assumed to be constant). Then we ran the subspace iteration and approximated matrix-vector products and linear combinations in appropriate tangent spaces (TS) of low-rank TT manifolds. This approach avoids rank increase and is therefore more efficient, e.g., with approximate matrix-vector products between $A$ and $v$ costing $\mathcal{O}(dnr_{v}^{3})$ where $Av$ is approximated in the tangent space of the low-rank manifold at $v$. fig. 7 shows the norm of the residual for each Ritz vector at each iteration. As the maximum truncation rank increases, the final accuracy of the Ritz vectors improves. Notably, different truncation ranks and truncation strategies (SVD or TS) do not significantly affect the convergence rate of the method, even when they influence the final accuracy. These observations highlight the robustness of the subspace iteration convergence rate to rank truncation errors. The $100$ iterations using TT-SVD truncation took 1896 seconds while $100$ iterations with tangent space projections took 656 seconds, more than 2.5x speed-up.
@@ 7 Conclusions
We examined low-rank inexact variants of the Lanczos method and a polynomial filtered subspace iteration for tensor eigenvalue problems. These methods typically avoid the issue of getting stuck at local minimizers, which can occasionally occur in DMRG-based algorithms. Furthermore, these methods can be easily used to compute more than one eigenpair of a general matrix, whereas the original DMRG algorithm is designed to compute the lowest eigenvalue (i.e., the ground state) of a Hermitian matrix. Our results demonstrated that orthogonal Krylov basis vectors often cannot be accurately represented by low-rank tensor trains, even when the eigenvectors of the associated matrix exhibit low-rank structure. We further showed that the low-rank approximation errors from representing orthogonal Krylov basis vectors resulted in stagnation of convergence in the Lanczos method. In contrast, the subspace iteration method employs Ritz vectors as the subspace basis, which directly approximate the low-rank eigenvectors of interest. We provided a convergence analysis quantifying the acceptable truncation error that ensures progress toward convergence at each iteration. Numerical experiments further confirmed that the subspace iteration method exhibits significantly greater robustness to truncation errors compared to the Lanczos method. We provided numerical demonstrations of the convergence analysis and showed that polynomial filtered subspace iteration can be used to compute several dominant eigenpairs with machine precision for problems with up to $100$ dimensions, provided the eigenvectors admit low-rank representations. We also demonstrated that rank-truncated subspace iteration can converge for problems where the density matrix renormalization group method (DMRG). Although the analysis presented here focuses on polynomial filtered subspace iteration, it can be easily extended to other type of subspace iteration such as the one based on the FEAST algorithm [2].
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program through the FASTMath Institute and Partnership with Basic Energy Science under U.S. Department of Energy Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAPm1027.
Code availability
Source code to reproduce numerical experiments in this paper are available at https://github.com/adektor/TT_subspace_eig.
Appendix A Lanczos method
$\mathcal{K}(A,v_{0};m)=\text{span}\{v_{0},Av_{0},...,A^{m-1}v_{0}\}.$ (49)
$AV_{m}=V_{m}T_{m}+fe_{m}^{\text{T}},$ (50)
$\beta_{j+1}v_{j+1}=Av_{j}-\alpha_{j}v_{j}-\beta_{j}v_{j-1},$ (51)
A.1 Lanczos with rank truncation
$\tilde{T}_{m}=R^{-\text{T}}\left(V_{m}AV_{m}\right)R^{-1},$ (52)
Algorithm 5 Lanczos iteration with rank truncation
1:
[…]
6:
7:  Approximate Krylov basis vectors $(v_{0},v_{1},\ldots,v_{m-1})$ in TT format
[…]
11:for $j=1$ to m1m-1 do
12:  $\beta_{j}\leftarrow\|v_{j}\|$ $\triangleright$ Compute subdiagonal element
13:  $v_{j}\leftarrow v_{j}/\beta_{j}$ $\triangleright$ Normalize
14:  $v_{j+1}\leftarrow\mathfrak{T}_{\bm{r}}(Av_{j})$ $\triangleright$ Apply AA and truncate
15:  $\alpha_{j}\leftarrow\langle v_{j+1},v_{j}\rangle$ $\triangleright$ Compute diagonal element
16:  $v_{j+1}\leftarrow\mathfrak{T}_{\bm{r}}(v_{j+1}-\alpha_{j}v_{j}-\beta_{j}v_{j-1})$ $\triangleright$ Orthogonalize and truncate
17:end for
A.2 Convergence of Lanczos with rank truncation
$V_{m}=\hat{V}_{m}+E_{m},$ (53)
Projecting $A$ into an orthonormal basis of $V_{m}$ yields a $m\times m$ projected matrix that can be viewed as a perturbation of the tridiagonal matrix $T_{m}$ in (50) generated by truncation-free Lanczos iteraion. However, since the perturbation can be large, the eigenvalues of such a projected matrix (the Ritz values relative to the inexact Krylov subspace) can be significantly different from the eigenvalues of $T_{m}$ (the Ritz values of the Krylov subspace). Moreover, a Ritz vector obtained from the approximate Lanczos procedure cannot be expressed as $f_{m-1}(A)v_{0}$ for some polynomial $f_{m-1}(\lambda)$ of degree m1m-1. Consequently, the accuracy of the approximation cannot be readily analyzed using the polynomial approximation theory typically employed to study the convergence of the Lanczos algorithm [27]. It is, however, well-known that large errors during the early iterations of the Lanczos method, such as those introduced by rank truncation, can lead to stagnation in convergence [25]. This stagnation is observed in our numerical experiments presented in section 6.
Appendix B Orthogonal projection onto TT tangent space
The collection of all TTs with rank-$\bm{r}$
$\mathcal{M}_{\bm{r}}=\{v\in\mathbb{C}^{n_{1}\times\cdots\times n_{d}}\mid\text{TT-rank}(v)=\bm{r}\},$ (54)
forms a smooth submanifold of $\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$ [13]. Associated with any $v\in\mathcal{M}_{\bm{r}}$ is a tangent space denoted by $T_{v}\mathcal{M}_{\bm{r}}$, which is a vector subspace of $\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$. Such tangent spaces, typically used for dynamical low-rank approximation of tensor differential equations [22, 11], are a convenient set of vector subspaces in which we can approximate the linear algebra operations needed for inexact subspace projection methods. Given the TT representation (3) of $v$, any element of the tangent space can be written (non-uniquely) as
$\delta v=\delta C_{1}C_{2}\cdots C_{d}+C_{1}\delta C_{2}C_{3}\cdots C_{d}+\cdots+C_{1}\cdots C_{d-1}\delta C_{d},$ (55)
Appendix C Proof of Theorem 4.1
Proof: Tangent of the angles $\measuredangle(v,\psi_{1})$ and $\measuredangle(Av,\psi_{1})$ can be expressed as
$\tan\measuredangle(v,\psi_{1})=\frac{\left\|\mathcal{P}_{\perp}v\right\|}{\left\|\mathcal{P}v\right\|}\quad\text{and}\quad\tan\measuredangle(Av,\psi_{1})=\frac{\left\|\mathcal{P}_{\perp}Av\right\|}{\left\|\mathcal{P}Av\right\|}.$ (56)
@@ Appendix B Orthogonal projection onto TT tangent space
$\left\|\mathcal{P}_{\perp}Av\right\|\leq|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|\quad\text{and}\quad\left\|\mathcal{P}Av\right\|=|\lambda_{1}|\left\|\mathcal{P}v\right\|.$ (57)
@@
$\left\|\mathcal{P}_{\perp}Av\right\|=\left\|\sum_{j=2}^{n_{1}\cdots n_{d}}\alpha_{j}\lambda_{j}\psi_{j}\right\|\\ \leq|\lambda_{2}|\left\|\sum_{j=2}^{n_{1}\cdots n_{d}}\alpha_{j}\psi_{j}\right\|\\ =|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|,$ (58)
@@
$\tan\measuredangle(Av,\psi_{1})=\frac{\left\|\mathcal{P}_{\perp}Av\right\|}{\left\|\mathcal{P}Av\right\|}\leq\frac{|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|}{|\lambda_{1}|\left\|\mathcal{P}v\right\|}\\ =\left|\frac{\lambda_{2}}{\lambda_{1}}\right|\tan\measuredangle(v,\psi_{1}),$ (59)
completing the proof. $\square$
Appendix D Proof of Theorem 4.3
Proof: Using the eigendecomposition $A=\Psi\Lambda\Psi^{\text{T}}$, we write $T_{AV}$ as
$T_{AV}$ $=\Psi_{>m}^{\text{T}}AV\left(\Psi_{\leq m}^{\text{T}}AV\right)^{-1}$ (60)
@@
$\tan\Theta(\mathcal{R}(AV),\mathcal{R}(\Psi_{\leq m}))$ $=\sigma_{1}(T_{AV})$ (61)